### v1

# Sub-GeV Dark Matter Direct Detection with Neutrino Observatories

Abstract
We present a new technique for sub-GeV dark matter (DM) searches and a new use of neutrino observatories. DM-electron scattering can excite or ionize target molecules in the observatory, which then produce light that can be detected by its photomultiplier tubes (PMTs). While individual DM scatterings are undetectable, the aggregate rate from many independent scatterings can be isolated from the total PMT dark rate using the expected DM annual modulation. We showcase this technique with the example of JUNO, a 20,000 ton scintillator detector, showing that its sensitivity in some mass ranges exceeds all other techniques and reaches key particle-theory targets.
[…]
It is very hard to test dark matter (DM) in the MeV–GeV mass range. A key challenge is the low recoil energies in scattering with matter. Compared with DM in the GeV mass range, the allowed cross sections are several orders of magnitude larger, making them tantalizing targets. Therefore, improving sensitivity to sub-GeV DM is a major goal of the particle physics community [1]. There has been a surge of interest in novel detection strategies, including using superconductors, superfluids, organic targets, quantum devices, and more [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 2, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 1, 37, 38, 39, 40, 41].
In this Letter, we present a new technique for sub-GeV DM searches and a new use of neutrino observatories. The basic ideas are as follows. DM-electron scattering within a neutrino observatory excites or ionizes target molecules, producing eV-range photons that can be detected with the photomultiplier tubes (PMTs). Individual DM-electron scattering events are undetectable because they typically lead to less than one PMT hit. However, we show that neutrino observatories can search for the aggregate signal rate from many independent DM scattering events. This signal rate must be compared to the observatory’s dark rate, which is the total rate of PMT hits due to thermal noise and detector radioactivities. Though the DM signal rate is small compared to the dark rate, it can be isolated using the annual modulation due to Earth’s orbital motion. This is a new use of neutrino observatories, which normally reconstruct signals and differentiate them from backgrounds using many correlated PMT hits. See Fig. 1.
We illustrate this new technique through the example of the Jiangmen Underground Neutrino Observatory (JUNO) [42, 43]. The first advantage is a huge target volume — here 20,000 ton — whereas direct-detection experiments have only reached the ton scale. The second is that the detector material — linear alkylbenzene (LAB), a liquid scintillator — has an electronic excitation threshold of only a few eV, whereas noble liquids have thresholds of order 10–100 eV. The obvious disadvantage of neutrino observatories compared to direct-detection experiments is much larger backgrounds (both absolute and relative). However, we show that the net effect of these factors is that JUNO has sensitivity to DM-electron interactions in some mass ranges that (A) exceeds all other techniques and (B) reaches some important theoretical targets. Although we focus on JUNO, our general ideas can be adapted to other neutrino observatories ($e.g.$, Super-K/Hyper-K, DUNE, and IceCube). This new use of neutrino observatories would complement familiar techniques, $e.g.$, searches for DM annihilation products [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61] or the scattering of boosted [62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77] or heavy or strongly interacting DM [78, 79, 80, 81, 82, 83, 84].
Figure 1: Schematic of our new technique. Multiple independent DM scatterings with molecules in a neutrino observatory produce photons that are detected by the PMTs.
In the following, we review the JUNO detector and DM scattering with bound electrons, then calculate the DM signal rate and the DM sensitivity when using annual modulation to suppress backgrounds. Finally, we conclude and discuss the prospects for further developments. In our calculations, we make several conservative assumptions, sometimes by potentially more than an order of magnitude. In forthcoming work, we will present more details on all aspects.
Review of the JUNO Experiment.— The JUNO experiment [42, 43], located in southern China, is planned to begin operations in 2025. The central goal of JUNO is to probe the neutrino mass ordering by precisely measuring the mixing of reactor antineutrinos. It will also make a wide variety of other important measurements in neutrino physics and astrophysics.
To achieve its goals, JUNO needs excellent light production and detection. The optically isolated inner volume of JUNO, which contains 20 kton of LAB, also contains other compounds to increase the production and transmission of light. There is a small admixture (2.5 g/L) of PPO (2,5-diphenyloxazole) as a fluor and a tiny admixture (3 mg/L) of bis-MSB (1,4-bis(2-methylstyryl)benzene), a wavelength shifter [85]. With this mixture, JUNO has high light production and transmission. On the enclosing sphere of the inner volume, 78% of the surface is covered with PMTs with high quantum efficiencies [86]. There are 17,612 20-inch PMTs and 25,600 3-inch PMTs; by detection area, the 20-inch PMTs dominate.
JUNO’s dark rate, dominated by thermal noise [87], will be high [86]. Its 20-inch PMTs each trigger at about 25 kHz, while its 3-inch PMTs each trigger at about 5 kHz, leading to a total dark rate rate of about $6\times 10^{5}$ kHz, $i.e.$, about $5\times 10^{13}$ PMT hits per day and $2\times 10^{16}$ per year. Those rates were obtained at surface temperatures and background levels, plus after only short periods in darkness. Once the PMTs are deployed in JUNO, all of those factors will be more favorable to lower rates, which would only help our results. For our purposes, it will be enough for JUNO to track the total number of hits per small time interval, which we take to be one day.
Review of DM-Electron Scattering.— Upon scattering with a molecule, DM can excite an electron from the ground state into one of several low-lying unoccupied orbitals that can de-excite by photon emission. In the medium, this de-excitation photon can be absorbed and re-emitted as a longer-wavelength fluorescence photon that can propagate through the volume and be detected by the PMTs.
As this process involves bound electrons, there are important dynamical effects [5, 24]. First, bound electrons have typical speeds about an order of magnitude faster than the average DM particle. Second, bound electrons have a distribution of speeds up to high values. These effects substantially change the scattering kinematics and break the simple link between momentum transfer and energy deposition used for nucleon scattering [5]. Third, the molecular energy levels have a complicated structure, requiring the use of a form factor $f_{ij}$, discussed below.
Still, with some simple arguments, we can understand some key scales. The energy transferred to the electron (of mass $m_{e}$), $\Delta{E_{e}}$, is related to the momentum lost by the dark matter (of mass $m_{\chi}$ and incoming velocity $\vec{v}$), $\vec{q}$, via energy conservation [5],
$\Delta{E_{e}}=-\Delta E_{\chi}=-\frac{|m_{\chi}\vec{v}-\vec{q}|^{2}}{2m_{\chi}}+\frac{1}{2}m_{\chi}v^{2}=\vec{q}\cdot\vec{v}-\frac{q^{2}}{2m_{\chi}}\,,$ (1)
where we have neglected the tiny molecular recoil. The maximum energy transfer is then [5]
$\Delta{E_{e}}\leq\frac{1}{2}m_{\chi}v^{2}\simeq\frac{1}{2}\,\text{eV}\times\left(\frac{m_{\chi}}{\text{MeV}}\right)\,.$ (2)
The high speeds of bound electrons relative to DM enhance the likelihood of excitations that are sufficient to overcome this threshold. For LAB with an excitation threshold of about 5 eV, this means that excitation is generally viable for DM with a mass greater than about 10 MeV.
DM Signal Rate.— We now calculate the aggregate DM signal rate in JUNO. The reaction rate factor, $\sigma v$, for DM particles to excite molecular electrons from level $i\rightarrow j$ (we assume $i$ is always the ground state) is [5]
$(\sigma v)_{ij}=\int\frac{d^{3}\vec{q}}{(2\pi)^{3}}(2\pi)\delta(\Delta E_{ij}-\omega)\frac{\absolutevalue{\mathcal{M}_{\text{free}}}^{2}\absolutevalue{f_{ij}(\vec{q})}^{2}}{16m_{\chi}^{2}m_{e}^{2}},$ (3)
where $\Delta E_{ij}$ is the difference in molecular energy levels, $\mathcal{M}_{\text{free}}$ is the amplitude for DM scattering on free electrons, and the energy transfer to the molecular target is $\omega=\Delta E_{\chi}$ per Eq. (1). There are two form factors, $F_{\text{DM}}$ and $f_{ij}$. The first parameterizes the momentum dependence of the amplitude and depends on the particle DM model, such that
|$\frac{\absolutevalue{\mathcal{M}_{\text{free}}}^{2}}{16m_{\chi}^{2}m_{e}^{2}}\equiv\frac{\pi\bar{\sigma}_{e}}{\mu_{\chi e}^{2}}F_{\text{DM}}^{2}(q),$ (4)
where $\mu_{\chi e}$ is the reduced mass of the DM-electron system and $\bar{\sigma}_{e}$ is the DM-electron scattering cross section, where the bar implies evaluation at the conventional fixed momentum transfer of $q=\alpha m_{e}$.
The second form factor, $f_{ij}$, parameterizes the molecular response to excitation. LAB is a class of organic compounds with the chemical formula C6H5CnH2n+1; in JUNO, n = 10–13 [88]. LAB contains a benzene ring, which features alternating double bonds, resulting in a system of conjugated $\pi$-bonds, which are covalent. The electrons occupying these $\pi$-orbitals are effectively delocalized, allowing them to move freely around the aromatic ring. In six-membered aromatic rings, six electrons participate in this conjugated $\pi$-electron system. Scintillation occurs due to electronic transitions where these $\pi$-electrons are excited into higher-energy molecular orbitals and then de-excite by photon emission.
To the best of our knowledge, LAB has not been considered as a sub-GeV DM detection material, but similar aromatic organic targets have been, $e.g.$, EJ-301 [22], which also has a benzene ring. In Ref. [22], the form factor was simply taken to be that for benzene, as the $\pi$-electrons in the benzene ring are expected to dominate the rate. While the lowest-energy transition is suppressed for benzene alone (its symmetry imposes selection rules that suppress some electronic transitions), it will not be for LAB, because the long alkyl chain attached to the benzene ring lowers the symmetry of the molecule, changing the form factor. To be conservative, we also use the benzene form factor alone [22]. We expect that a proper treatment of the LAB form factor will improve sensitivity. We will treat this in detail in future work.
To obtain the total DM scattering rate, $R_{\rm scatter}$, we include the DM velocity distribution, $g_{\chi}(\vec{v})$,
$R_{\rm scatter}=\sum_{i,j}N^{e}_{\rm LAB}\frac{\rho_{\chi}}{m_{\chi}}\int d^{3}\vec{v}g_{\chi}(\vec{v})(\sigma v)_{ij}\,,$ (5)
where $\rho_{\chi}=0.4$ GeV/cm3 is the local DM mass density, and $N^{e}_{\rm LAB}\simeq 3\times 10^{32}$ is the number of LAB molecules in JUNO’s inner volume multiplied by the six $\pi$-electrons in LAB’s benzene ring. For the total rate, we approximate $g_{\chi}(\vec{v})$ to be spherically symmetric, assuming a Maxwell-Boltzmann distribution with a cutoff at the escape velocity of the Milky Way,
$g_{\chi}(v)=\frac{1}{K}\exp\left(-\frac{\absolutevalue{\vec{v}+\vec{v_{E}}}^{2}}{v_{0}^{2}}\right)\Theta(v_{\text{esc}}-\absolutevalue{\vec{v}+\vec{v_{E}}})\,,$ (6)
[…]
$\eta(v^{ij}_{\text{min}})$ $=\int\frac{4\pi v^{2}dv}{v}g_{\chi}(v)\Theta(v-v^{ij}_{\text{min}}(q))\,,$ (7)
$v^{ij}_{\text{min}}(q)$ $=\frac{\Delta E_{ij}}{q}+\frac{q}{2m_{\chi}}\,,$ (8)
where $\eta(v^{ij}_{\text{min}})$ is the integrated velocity distribution of DM particles above the threshold $v_{\rm min}^{ij}(q)$, which is the minimum speed DM needs for the electron to gain an energy $\Delta E_{ij}$ with momentum transfer $q$.
The expected signal rate in JUNO is therefore
$R_{\rm sig}=\xi\frac{N^{e}_{\rm LAB}\rho_{\chi}\bar{\sigma}_{e}}{8\pi m_{\chi}\mu_{\chi e}^{2}}\sum_{i,j}\int\frac{d^{3}\vec{q}}{q}\eta(v^{ij}_{\text{min}}(q))F_{\text{DM}}^{2}(q)\absolutevalue{f_{ij}(\vec{q})}^{2},$ (9)
where we have multiplied the DM scattering rate, $R_{\rm scatter}$, by an efficiency factor $\xi$, which we find to be $\xi=2.2\times 10^{-2}$. We obtain this as
$\xi=Q\times X\times Y\,,$ (10)
where $Q$ is the fluorescence quantum yield of LAB [89], $X$ is the wall coverage of the PMTs, and $Y$ is the quantum efficiency of the PMT photocathode integrated over the emission spectrum of JUNO’s active medium, taking into account absorption and emission of photons from the primary scintillating target by the fluor and wavelength shifter. For JUNO, we use $Q=0.1$, $X=0.78$, and $Y=0.28$ [89, 90, 91, 92]. In Eq. (9), the key other input that substantially changes the rate compared to a naive estimate (based on the product of the DM number density, velocity, number of targets, and cross section) is the effect of the molecular form factor $f_{ij}$, which increases the rate by a few orders of magnitude.
In Eq. (9), we conservatively assume the rate purely arises from excitation and fluorescence of the LAB (and consequently the fluors/wavelength shifters). In forthcoming work, we will consider additional signals from ionization (where the electron is ejected entirely, and the molecule later undergoes recombination) and propagation (where an ionized electron can excite or ionize other molecules through collisions).
DM Sensitivity from Annual Modulation.— In many dedicated DM detectors, the backgrounds are small after cuts, $e.g.$, requiring that events arise within a fiducial volume, occur in a target energy range, or have certain characteristics. Here, such cuts are not possible because the only observable is the total rate of all PMT hits. This rate is dominated by the background (the dark rate, $B\sim 2\times 10^{16}$ events/yr). Because the background cannot be predicted precisely enough, we cannot estimate the DM signal sensitivity by requiring $S_{\rm tot}\gtrsim\sqrt{B_{\rm tot}}$, where those are the total numbers of signal and background events in the observation period, $T$.
However, we can isolate even a small DM signal through the annual modulation caused by Earth’s orbital motion [93], which changes the DM flux that appears via Eq. (9). This rate is used on 2 March and 2 September, while the DM average velocity increases by 15 km/s on 2 June and decreases by 15 km/s on 2 December. This leads to a harmonic modulation amplitude $f_{\text{mod}}$, which is approximately 0.1 for most DM masses; it is somewhat larger (smaller) for light (heavy) DM masses. There is no trials factor in searching for this modulation, as its amplitude, period, and phase are all known.
[…]
To illustrate how to isolate a DM signal, we first consider the simplest case, where the background rate is constant, $B(t)=B_{\text{tot}}/T$. The DM signal is
$S(t)=\frac{S_{\text{tot}}}{T}\left[(1+f_{\text{mod}}\cos(\frac{2\pi t}{\text{1yr}}-\frac{\pi}{2})\right]\,.$ (11)
For convenience here, we assume that $T$ is an integer number of years, with the time interval starting on 2 March, so that the DM signal will be increased during the first half year of data collection. The amplitude of the varying part of the signal, $S_{\text{tot}}f_{\text{mod}}/T$, can be extracted by fitting the binned time-series data for $N(t)=B(t)+S(t)$. The basic results and uncertainties can also be obtained via an asymmetry analysis [102, 103], using
$A=\frac{N_{\text{inc}}-N_{\text{dec}}}{N_{\text{inc}}+N_{\text{dec}}}=\frac{S_{\text{tot}}f_{\text{mod}}}{B_{\text{tot}}}\,,$ (12)
where $N_{\text{inc}}$ is the number of events integrated over the first half of the year, when the DM signal is increased, and $N_{\text{dec}}$ is its complement. When the asymmetry is small, as it is here, the uncertainty on $A$ is
$\delta A\simeq\frac{1}{\sqrt{N_{\text{tot}}}}\simeq\frac{1}{\sqrt{B_{\text{tot}}}}\,.$ (13)
Requiring $\delta A/A\gtrsim 1$ for signal detection leads to requiring $S_{\text{tot}}\gtrsim\sqrt{B_{\text{tot}}}/f_{\text{mod}}\sim 10\sqrt{B_{\text{tot}}}$.
Measured dark rates, $e.g.$, by Super-K [104] and IceCube [105], are not quite constant. There are non-statistical fluctuations over short timescales, but these will integrate out in the calculation of the asymmetry. There are also slow, secular trends in the dark rate. Over a timescale of one year, trends like these could introduce asymmetries, but shifting the analysis window by half a year would show that the asymmetries are spurious, because a DM signal asymmetry would change sign. With years of data, these secular trends could be fit and removed. Of greater concern are components of the background that have or could have annual modulations, some of which are discussed in Ref. [106]. PMT thermal noise depends on the detector temperature, radon and muons depend on atmospheric conditions, solar neutrinos depend on the distance to the Sun, and reactor antineutrinos depend on power cycles. Importantly, the rates of those backgrounds can be precisely monitored via independent data. Even if they could not, they are shifted in phase from the DM signal by weeks to months and will have small phase uncertainties. Finally, we note that combining the results of multiple detectors would make the analysis even more robust, as the signals on the properties of the detector, and backgrounds depend on that and the detector location.
Figure 2 shows the projected sensitivity to the DM-electron scattering cross section as a function of DM mass, for a 95% C.L. and one year of JUNO data. We show results for two benchmark scenarios for the DM-electron interaction: via a heavy mediator (so $F_{\text{DM}}=1$) and via a light mediator (so $F_{\text{DM}}(q)=(\alpha m_{e})^{2}/q^{2}$). We also show complementary constraints from direct-detection experiments and solar reflection. The orange shaded regions show the community’s theory target regions for benchmark freeze-in models that produce the observed relic DM abundance.
The position of the minimum near 10 MeV in Fig. 2 is driven by the low-excitation threshold of LAB being a few eV, as discussed earlier. At higher masses, the sensitivity degrades linearly due to the lower DM number density. At lower masses, it degrades rapidly because fewer DM particles have sufficient momentum to overcome the excitation threshold. Further, the modulation fraction grows as the DM becomes lighter. This changes the shape of the curve, especially for DM masses near the threshold, such that the minimum is moved to slightly lower DM masses (in agreement with the modulation results of, $e.g.$, Ref. [5]).
Conclusions and Prospects.— The nature of DM remains one of the most important open questions in physics. Despite extensive efforts, a wide array of DM experiments have yet to yield definitive signals. This has motivated increasing interest in exploring broader possibilities than GeV-range mass WIMPs and other theoretically motivated DM candidates. For example, DM candidates in the sub-GeV mass range are the subject of intense focus now. However, detecting sub-GeV DM is notoriously difficult to detect, due to the small energy transfers in their scattering interactions.
We introduce a new technique to probe sub-GeV DM and a new use of neutrino observatories. Specifically, we demonstrate that DM-electron scattering in neutrino observatories can produce detectable signals through excitation of target molecules. While individual scattering events are undetectable, the aggregate rate from many independent scatters can be isolated from the PMT dark rate through annual modulation. Focusing on JUNO, we show that this method enables world-leading sensitivity to the DM-electron scattering cross section in some mass ranges, reaching cross sections as low as a few times $10^{-39}$ cm2 for MeV-scale DM with one year of data. This power comes from JUNO’s huge size and high fluorescence quantum yield.
Next steps include refinement of the theoretical and experimental details for this technique in JUNO. This requires a deeper understanding of the detector’s background rates, improved modeling of the signal response, and a detailed study of potential systematic uncertainties. With JUNO’s data taking starting in 2025, early-stage experimental validation of key inputs would accelerate the process. While there will be experimental challenges in implementing our new technique, we are optimistic that they can be surmounted, given the importance of the task.
Beyond JUNO, our idea can be applied to other neutrino observatories, such as Super-K/Hyper-K, DUNE, and IceCube. While these detectors have less light emission than JUNO, they also have lower dark rates. Utilizing multiple detectors would serve two key purposes. First, making searches more robust by comparing different detectors, as they should observe consistent amplitudes and phases for an annual modulation signal, which does not depend on detector location, though some backgrounds do. Second, extracting more detailed DM properties. If a signal were observed in JUNO, we would extract only a degenerate combination of DM mass and cross section, which would appear as a characteristic trajectory in Fig. 2. Independent observations could help disentangle the specific DM mass and cross section, as the trajectories depend on the detector properties. It would also be interesting to explore the potential of this technique for dedicated DM detectors. We will explore these topics in future work.
Ultimately, this work opens a new frontier in DM detection. Neutrino observatories, which are designed for completely different physics, may hold the key to uncovering sub-GeV DM. The potential is now clear, but realizing it will require dedicated experimental validation, sharper theoretical modeling, and coordinated multi-detector searches. If DM hides in this mass range, we may already have the tools to find it. We just need to start looking, but in a new way.
Acknowledgments.— We are grateful for helpful discussions with Carlos Blanco, Yufeng Li, Bernhard Mistlberger, Mark Vagins, and Tien-Tien Yu. RKL was supported by the U.S. Department of Energy under Contract DE-AC02-76SF00515. JFB was supported by National Science Foundation Grant No. PHY-2310018.